### v1

# Towards Interactive Autonomous Vehicle Testing: Vehicle-Under-Test-Centered Traffic Simulation thanks: This work was supported in part by National Natural Science Foundation of China 52125208, 52232015

Abstract
The simulation-based testing is essential for safely implementing autonomous vehicles (AVs) on roads, necessitating simulated traffic environments that dynamically interact with the Vehicle Under Test (VUT). This study introduces a VUT-Centered environmental Dynamics Inference (VCDI) model for realistic, interactive, and diverse background traffic simulation. VCDI is built on a Transformer-based trajectory inference model to generate trajectories for background objects. Serving the purpose of AV testing, VCDI additionally considers VUT-centered interactivity and scenario diversity using a conditional inference framework. First, the VUT future motion is taken as an augmented model input to bridge the interaction between VUT and background objects. Second, to enrich the scenario diversity, a Bayesian-network-based cost function module is designed. The module, learned in a distributional form, captures the uncertainty of the VUT’s strategy, triggering various scenario evolution. Experimental results validate VCDI’s trajectory-level simulation precision which outperforms the state-of-the-art trajectory prediction work. The flexibility of the distributional cost function allows VCDI to provide diverse-yet-realistic scenarios for AV testing. We demonstrate such capability by modifying the anticipation to VUT’s cost-based strategy and thus achieve multiple testing scenarios with explainable background traffic evolution.
Index Terms:
autonomous vehicles, simulation-based testing, background traffic simulation, driving interaction
I Introduction
The deployment of autonomous vehicles (AV) on public roads is forming a human-machine mixed traffic environment. AVs, as new traffic participants, are found to possibly compromise road safety when interacting with human road users [1, 2]. Waymo’s 2020 report revealed that 78% of AV-involved accidents resulted from misunderstandings of AV’s actions [3], indicating AV’s insufficient interacting ability. The human-machine mixed driving environment poses new challenges for the interactive capabilities of autonomous driving and also brings new requirements for testing these systems.
AV testing is fundamentally tied to realistic, interactive, and diverse simulation environments where background dynamic participants interact with the vehicle under test (VUT), namely the AV. In pursuing the driving behaviors of background traffic participants that are consistent with real-world traffic, efforts have been widely rolled out on the following challenges, realism of the scene setting, inter-agent interactivity, and diversity of scenario evolution. First, the realism of the scene setting ensures that the testing scenario is possibly encountered in real-world traffic, thus providing testing value. Second, inter-agent interaction calls for congruent action-dependence among VUT and background traffic participants. Third, the diversity of scenario evolution is essential as it reflects the inherent uncertainty and dynamics of real-world traffic, ensuring that the simulation environment authentically represents the wide range of situations VUT might encounter during normal operations.
Figure 1: VUT-Centered Environmental Dynamics Inference Framework.
Focusing on varied points, solutions have been developed to generate testing environments in the simulation while handling the above challenges. The log test is primarily applied to ensure realism by directly replaying the real-world trajectories performed by human drivers while suffering from the lack of interaction with VUT [4]. By introducing behavior models like the Intelligent Driver Model (IDM) to enable background participants’ responses to VUT [5], the model-based test advances the testing environments to be dynamic. While a step towards interactivity, the model-based test often results in overly cautious background object behaviors, avoiding collisions and failing to replicate high-risk scenarios that are crucial for testing. The learning-based test generates interaction-aware background participants by learning implicit interactive relationships from driving data [6]. However, the interactivity does not extend to VUT which means background participants take VUT as a non-strategic moving obstacle, isolating VUT from the multi-agent interaction.
Serving the purpose of trustworthy AV testing, this work proposes a VUT-Centered environmental Dynamics Inference (VCDI) framework for simulating background traffic participants that interact with VUT in a human-like manner. Specifically, VCDI applies a Transformer-based neural network model to infer the future trajectories of traffic participants (Fig. 1). We initialize the simulation with real-world driving scenario records to ensure the realism of the scenario setting. To capture background traffic objects’ responses to the VUT, the anticipated trajectory of VUT is taken as additional model input, facilitating VUT-centered interactivity. For the application of AV testing where VUT’s strategy is unavailable, the strategy presets of VUT are generated by a distributional cost function, which is learned from real-world AV-involved driving data. By varying the strategy presets, VCDI could simulate diverse-yet-realistic background traffic dynamics from real-world initial scenario states.
The contributions of this work are three-fold:
  • A conditional trajectory inference framework VCDI is proposed to model background traffic objects’ strategic responses, centering simulated environmental dynamics on the VUT.
  • A distributional cost function module is designed and integrated to provide flexible and explainable inference conditions, enabling diverse traffic simulation.
  • VCDI’s potential for directional background traffic simulation is demonstrated through experiments in typical traffic scenes.
The remainder of this work starts from a review of related works (Section II) which systematically analyzes the pros and cons of background traffic simulation techniques for AV testing from previous efforts. We then introduce the design details of VCDI framework in Section III and the validation settings in Section IV. Results are discussed in Section V to demonstrate VCDI’s capability of simulating diverse traffic environments. Finally, we conclude our work in Section VI.
II Related Works
The pursuit of autonomous driving safety through on-road testing presents substantial challenges and costs. It has been estimated that billions of kilometers are required to validate the safety of AV, an endeavor that is impractical in terms of both time and resources [7]. Consequently, the focus has shifted towards simulation-based methodologies where challenges lie in the reproduction of real-world traffic that interacts properly with VUT [4].
The log test, which replays recorded trajectory data from actual driving scenarios, offers high reproducibility, enabling researchers to consistently test VUT under controlled conditions [8, 9, 10]. As an extension, Wang et al. extract principle components of the traffic, termed traffic primitives, from the logged data to guide new scenario generation, enriching the log test from being log-only [11]. However, the static nature of the log test is a major drawback. Since all scenarios are predetermined, the log test fails to account for the dynamic nature of real-world driving where traffic conditions are constantly affected by VUT.
To address the non-interactive limitations of the log test, analytical models [12, 13] like the Intelligent Driver Model (IDM) were applied [14], constructing the model-based test [15]. These models allow background objects to respond to VUT’s actions, providing a certain level of interaction. However, despite the introduction of interactivity, analytical models often result in overly cautious behavior from background vehicles, thus failing to simulate hazardous interaction patterns. The behaviors of background traffic participants in model-based test are still bounded by the model parameters and typically fail to replicate high-risk scenarios that are crucial for exhaustive testing.
Following the identification of limitations inherent in both the log test and the model-based test, the research community has sought integrated approaches that combine the inherent realism of the log test with the interactivity of the model-based test. The primary challenge lies in accurately replicating complex and multi-objective scenarios that not only demand a high degree of realism but also necessitate intricate interactions and strategic considerations among traffic participants.
Deep learning (DL) models emerge as a viable methodology in addressing these challenges, which can incorporate implicit interactive relationships among traffic participants [4, 16, 17, 18]. The key advantage of leveraging such implicit interactions lies in their abilities to simulate a more realistic traffic environment, where background vehicles can exhibit complex and human-like driving behaviors. Mütsch et al. have noted an emerging trend where DL models are progressively supplanting traditional model-based simulation in the fields of autonomous driving testing [19]. DL-based simulation models, despite their capabilities of capturing complex traffic patterns, primarily focus on the interaction among modeled agents, i.e., the background objects while isolating the VUT. Therefore, the effectiveness of the DL-based test is confined to reacting to observable VUT states, lacking the foresight to interpret and adapt to VUT’s underlying strategic objectives.
The progression from log test to model-based test and to DL-based test reflects the ongoing need for AV testing scenarios that are both realistic and interactive. In this work, we incorporate the strategy presets of VUT as inputs that influence the future trajectory inference of background participants. Concurrently, the strategy presets of VUT act as the keys to initiating scenarios. Different presets for VUT unlock varied scenario evolution, thereby providing a dynamic and adaptable testing environment.
III Method
III-A Problem Description
Our work is based on a DL-based trajectory prediction model [20] where trajectories of background traffic participants are inferred over a forecast horizon $T_{f}$. In addition to the historical states of all agents and environmental context, we further applied the anticipated future trajectory of VUT as input to capture background traffic participants’ responses to the strategy of VUT. Let $S_{VUT}^{fut}$ represent the future trajectory of VUT, which is a sequence of anticipated states over the forecast horizon. The input $X$ to the inference model includes, along with the current environmental information $M$, the historical states over a time period $T_{h}$ of VUT $S_{0}$, the historical states of $N$ surrounding agents $S_{1},S_{2}…S_{N}$, and the future trajectory of VUT $S_{VUT}^{fut}$. This enhancement aims to capture the reactive and strategic adjustments of surrounding agents to VUT’s future motion. The output of the trajectory inference model $Y$ is a trajectory set consisting of a sequence of 2D coordinates denoting the possible future position of the background traffic participants. The problem is formally expressed as:
$X$ $={M,S_{0},S_{1},S_{2}…S_{N},S_{VUT}^{fut}}$ (1)
$Y$ =$=\left\{\left(x_{i}^{t},y_{i}^{t}\right)\mid t\in\left\{t_{0}+1,\ldots,t_{0}+T_{f}\right\}\right\}_{i=1}^{N}$
where $S_{i}=\left\{s_{i}^{t_{0}-T_{h}+1},s_{i}^{t_{0}-T_{h}+2},\ldots,s_{i}^{t_{0}}\right\}$ is the dynamic state of the agent $i$ at timestep $t$, $(x_{i}^{t},y_{i}^{t})$ is the inferred coordinate of the target agent $i$ at timestep $t$, and $t_{0}$ is the current time step [21].
III-B VUT-Centered Environmental Dynamics Inference
As shown in Fig. 1, VCDI incorporates a Transformer-based neural network model for trajectory inference. The model processes the static input of the map through a Multilayer Perceptron (MLP), encoding the geometric and infrastructural context, and the dynamic input comprising the node history via a Long Short-Term Memory (LSTM) network, encapsulating the participants’ movement patterns. Additionally, the future trajectory of VUT is encoded using a LSTM network. These processed inputs are fed into the Transformer network, which captures spatial-temporal relationships between the elements within the traffic scene. Detailed specifications of the framework are outlined below.
1) Infer the trajectories of background traffic participants
Input Representation: Each agent’s history is captured as a sequence of dynamic features over the past 20 timesteps ($T_{h}$=20), including 2-D position, heading angle, velocity, and bounding box size. We focus on the nearest 10 agents ($N$=10) to VUT, storing their observations in a fixed-size tensor, padding absent agents with zeros. Scene context is derived from lanes and crosswalks, represented by vectors of waypoints with varying features. For each agent, a local scene context is constructed, highlighting probable lanes and nearby crosswalks. Lane waypoints feature center and boundary positions, headings, speed limits, and traffic signals. Crosswalk waypoints detail the perimeters of the crosswalk areas. Positional attributes of all agents and map elements are recalibrated to the VUT’s local coordinate frame. In addition, the VUT’s future trajectory is encapsulated as a dynamic feature sequence projected across the forthcoming 50 timesteps ($T_{f}$=50).
Encoding Scene Context: To encode historical states, the scene context is processed using a MLP. This involves encoding nearby lanes and crosswalks into vectors. These vectors represent numerical features such as positions and directions. The MLP enables discrete feature encoding and is essential for capturing the nuanced spatial relationships that influence an agent’s behavior.
Encoding Historical States and the Future Trajectory of VUT: Our model adopts a LSTM network to encode the historical states of agents, optimizing for predictive performance and computational efficiency, particularly in scenarios involving short-term time series data. To refine the inference accuracy and enhance temporal consistency, the framework employs a LSTM-Transformer hybrid model that is specifically tailored for encoding VUT’s future trajectory. This model integrates the LSTM’s capability to understand and remember long-term dependencies with the Transformer’s proficiency in handling sequences, resulting in a powerful representation of VUT’s dynamic behavior over time.
Interaction Graphs and Attention Mechanisms: We further refine the model by employing three interaction graphs: an agent-scene interaction graph, an agent-agent interaction graph and an action-action interaction graph. These graphs facilitate the understanding of complex dynamics and dependencies within the scene. Utilizing a two-layer self-attention Transformer encoder, the model can capture the intricate interaction relationships. This self-attention mechanism, which is inspired by the Transformer’s encoder, processes the encoded agents’ trajectory features and the map features to elucidate potential influences and interactions.
Decoding Futures: Decoding phase involves translating the interaction encodings into multimodal future trajectories for each agent. This is achieved by concatenating the agent-agent and agent-map interaction encodings, and processing them through a MLP to predict potential future states. The model is designed to output the trajectory with the highest probability for each agent as the final inferred trajectory.
2) VUT-centered augmented input
The unique setup of our model, as illustrated in Fig. 1, is the anticipated behaviors of VUT. Specifically, during the training phase, our model utilizes the actual future trajectory sequence of VUT as an additional input. This inclusion enables the model to infer the future trajectories of the surrounding traffic participants as they respond to VUT’s strategy. When simulating traffic environments based on the model, instead of using the actual trajectory, we input a planned trajectory by solving a cost-based optimization to handle the inaccessibility of the actual VUT’s future trajectory. The planned trajectory is generated based on the cost function learned during training. By sampling different VUT’s planned trajectories from the cost function, our model is capable of inferring the corresponding reactive trajectories of the background vehicles. This allows us to examine how surrounding traffic dynamically adjusts to various strategic presets of VUT, thus providing insights into the interactions between VUT and other road users within the traffic system.
3) Distributional cost function
The cost function of the model encompasses various essential driving aspects vital to VUT, such as speed, acceleration, jerk, steering angle, steering change rate and traffic rule adherence, focusing on staying near the lane centerline and following the lane direction. To promote efficiency, the function incorporates incentives that encourage VUT to maintain speeds within legal limits. Comfort is quantified through metrics evaluating acceleration, jerk, and steering. Moreover, compliance with traffic regulations is secured by imposing penalties for deviations from ideal lane positions and vehicular orientation. The parameters of the cost function is fine-tuned to minimize the divergence between actual trajectories from real-world data and the planning trajectories. The specific design references the methodology outlined in [20].
We utilize Bayesian networks to conceptualize the cost function for VUT as a distribution rather than being single-valued, reflecting the inherent uncertainty of traffic environments. To model and learn the distributional weight parameters of the cost function effectively, we assume that the cost function parameters follow a Gaussian distribution, which allows us to capture the essential characteristics of driving behavior with mean and variance as the primary descriptors. The Bayesian network updates the beliefs about the Gaussian parameters using Bayesian inference, which adjusts the mean and variance of the distribution accordingly [22].
III-C Learning process
In constructing the training process for our model, we adopted the loss function framework as described in the referenced work [20], adhering to an end-to-end methodology utilizing real-world driving data. The loss function is a composite of four terms, integrating prediction loss for all agents, score loss to calculate the probabilities of different futures, imitation loss, and the cost of planning for VUT. The comprehensive loss equation is defined as:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{prediction}}+\lambda_{2}\mathcal{L}_{\text{score}}+\lambda_{3}\mathcal{L}_{\text{imitation}}+\lambda_{4}\mathcal{L}_{\text{cost}}$ (2)
where $\lambda_{1}$ through $\lambda_{4}$ represent scaling weights assigned to these discrete loss components.
IV Experiments
IV-A Dataset
Our research employs the Waymo Open Motion Dataset for training and validating of our framework. This extensive dataset is instrumental in providing real-world insights, as it provides plenty of urban driving scenarios characterized by diverse road structures and varied traffic interactions. To dissect a 20-second driving scene from raw data into discrete frames, we establish a 7-second time window, inclusive of a 2-second historical observation span and a subsequent 5-second forecast horizon, sampled at 10 Hz. This window advances through the scene in increments of 10 timesteps (equivalent to 1 second), thereby generating 14 individual segments of 7-second frames.
For the model development, 80% of these selected scenes are designated for training, constituting a total of 35,968 scenarios. Another 10%, amounting to 4,498 scenarios, are used for validation, while the remaining 10% are reserved for testing.
IV-B Evaluation
We evaluate the performance of our model with two key metrics: the Average Displacement Error (ADE) and the Final Displacement Error (FDE) over three timesteps (1, 3, and 5 seconds). The metrics are computed for the prediction error of background objects to assess the model’s trajectory-level precision.
IV-C Ablation study
An ablation study is conducted to illustrate the specific model components that influence performance. Compared with the traditional DL-based trajectory prediction methods [23, 24, 20], we consider VUT’s strategy as an augmented input, which can result in background traffic participants adequately responding to VUT’s strategy. To reveal the importance of VUT-centered augmented input, we set up the work in [20] as a baseline model that drops out the augmented input. As shown in Table I, overall, the VCDI outperforms its baseline in the trajectory prediction tasks while a performance reduction is observed after the introduction of the distributional feature.
As the distributional cost function comes at the cost of raising unnecessary uncertainty from the view of trajectory prediction, the price of introducing Bayesian networks is studied. The single-valued version of our model, VCDI s.v., assumes a singular strategic behavior of VUT and infers a fixed VUT’s cost function. As shown in the Table I, the introduction of variability in strategic presets, although it leads to a reduction in aggregate inference precision, is a deliberate design choice that enhances the model’s capability to represent the diverse and uncertain nature of real-world driving behaviors. The case study we present later justifies this trade-off, showing that the decreased performance is offset by the benefits of accommodating diverse scenario evolutions.
TABLE I: ablation study in trajectory inference task
ADE FDE (1s) FDE (3s) FDE (5s)
No Augment 0.8624 0.2660 0.9631 2.0429
VCDI s.v. 0.8538 0.2529 0.9919 2.0255
VCDI 0.8612 0.2643 0.9620 2.0372
Figure 2: Diverse Scenario Evolution Guided by Distributional Cost Function
Figure 3: Car-following Scenario
Figure 4: Intersection Scenario
V Results and Discussion
V-A Distributional Cost Function
Fig. 2 illustrates the weights of the distributional cost function generated by our trajectory inference framework. A specific cost function, described by a set of weights from the distribution, serves as a crucial ”key” for activating scenarios. We could infer scenario evolution with keys either by sampling from the distribution or by making targeted selections.
Specifically, each key, whether sampled or selectively chosen, represents a strategic preset on VUT made by the background objects. The inference results then depict the future trajectories of background objects based on such presets, which allows for simulation that explicitly considers VUT’s strategy. In other words, given an initial scenario state extracted from real-world data, the inferred trajectories of background participants change according to the keys used for scenario activation.
V-B Typical Scenario Cases
The distributional cost function obtained from Section V-A serves as the key to initiating our testing scenarios. From this distribution, we select two different parameter combinations to construct the cost functions, which represent distinct presets of the background traffic participants towards VUT. Subsequently, simulation experiments are conducted in typical traffic scenes extracted from real-world driving data, including the car-following and intersection scenarios.
1) Car-following scenario: Fig. 3 illustrates both observed and inferred trajectory data. The leftmost subfigure shows actual data from a car-following scenario, while the two subfigures on the right present the trajectories of background participants, inferred based on various strategic presets applied to VUT. Specifically, we increased the emphasis on efficiency in VUT’s cost function to varying degrees. As a result of this intensified focus on efficiency, VUT is anticipated to pursue more travel progress, which in turn allows the following car to develop more efficiency-oriented trajectories.
2) Intersection scenario: Fig. 4 shows data from an intersection scene, where we also incrementally increased VUT’s preference for efficiency. It is evident that when VUT heightened its emphasis on efficiency, there was a noticeable shift in the semantic outcomes. Specifically, as the VUT’s focus on efficiency intensified, the behavior of the background vehicle changed from rushing to yielding to the VUT.
The results of the case study indicate that the different scenario activation triggers varied responses of the background participants. Furthermore, the evolution of simulation scenarios is to some extent controllable, as the key to scenario activation—the distributional cost function—is fully explainable.
VI Conclusion
The on-road driving safety of autonomous vehicles requires testing under scenarios that closely mirror real-world traffic conditions. In this paper, we propose a Vehicle Under Test (VUT)-Centered environmental Dynamics Inference (VCDI) framework to simulate background traffic objects. The VCDI’s unique capability to integrate a distributional cost function learned from real-world driving data enables the modeling of diverse, strategic responses from surrounding traffic participants to VUT’s maneuvers.
Our findings demonstrate that the VUT-centered augmented input enables background traffic participants driven by DL-based models to dynamically interact with VUT. In addition, the cost-based strategy presets on VUT support directional background traffic simulation, meriting the application of AV testing.